OpenStudy (anonymous):

I need help solving for X 4x + 5 = 25 and 7x/2+5=8

OpenStudy (anonymous):

4x + 5 = 25 4x = 25 - 5 4x = 20 x = 5

OpenStudy (goformit100):

equate both

OpenStudy (anonymous):

equate?

OpenStudy (anonymous):

And..so I'm just subtracting the bigger number with the smaller number? Also, the answers given only have X at the end, no numbers before it.

OpenStudy (anonymous):

Nvm. I saw how you did that with the 4. Sorry, I confused myself

OpenStudy (goformit100):

you r correct